OpenStudy (sbentleyaz):

The position equation for a particle is s(t)= sqrt{t ^{3}+1} where s is measured in feet and t is measured in seconds. Find the acceleration of the particle at 2 seconds the correct answer is 2/3 ft/sec2 but i do not know how you get to this answer

OpenStudy (sbentleyaz):

@3mar

OpenStudy (danjs):

recall for position s, velocity v, and acceleration a, at a time t... s(t) \[v(t)=s'(t)\] \[a(t)=v'(t)=s''(t)\]

OpenStudy (danjs):

instantaneous velocity is derivative of position, and instantaneous acceleration is derivative of velocity

OpenStudy (sbentleyaz):

Oh so just derive it twice and plug in 2? @DanJS

OpenStudy (sbentleyaz):

Right?

OpenStudy (danjs):

yes, to get the acceleration function, you take second derivative of position..

OpenStudy (sbentleyaz):

but how do you derive with the square root

OpenStudy (danjs):

\[\large \sqrt{u}=u^{1/2}\] \[\large s(t)=(t^3+1)^{\frac{ 1 }{ 2 }}\]

OpenStudy (sooobored):

chain rule and \[\sqrt{x} = x^{1/2}\] \[\frac{ d}{dx} x^n = n x^{n-1}\]

OpenStudy (danjs):

use the chain rule, for a quantity u, \[\large [u]^a = a*[u^{a-1}]*u'\]

OpenStudy (sbentleyaz):

so 1/2(t^3+1^-.5)(2t^2 ??

OpenStudy (sbentleyaz):

@3mar could you check if i derived that correctly? for the first derivative?

OpenStudy (danjs):

\[\large v(t)=s'(t)=\frac{ 3t^2 }{ 2\sqrt{t^3+1} }\]

OpenStudy (sbentleyaz):

Oh, thats what i meant. So then how would you do the second derivative

OpenStudy (sbentleyaz):

because we have another square root on the bottom

OpenStudy (sbentleyaz):

@3mar

OpenStudy (danjs):

the quotient rule works.. \[\frac{ d }{ dt }[\frac{ u }{ v }]=\frac{ v*u'-u*v' }{ v^2 }\]

OpenStudy (sbentleyaz):

ok thank you

OpenStudy (3mar):